Integrate tan³x dx using reduction method
1
Expert's answer
2020-02-20T09:15:21-0500

"I=\int \:tan^3\left(x\right)dx"

"=\int \:tan^2\left(x\right)\cdot tan(x)\:dx"

We know "tan^2(x)=sec^2(x)-1" . Substituting this in above integral.

"=\int \:\left(sec^2x-1\right)tan\left(x\right)dx"

"=\int \:\left(\:tan\left(x\right)sec^2\left(x\right)-\:tan\left(x\right)\right)dx"

"=\int \:tan\left(x\right)sec^2\left(x\right)dx-\int \:tan\left(x\right)dx"

"=I_1-I_2" ......................................................................................................eqn (1)

"I_1=\int tan(x)sec^2(x)dx"

apply u-substitution

Let "u=tanx"

"\Rightarrow du=sec^2(x)dx"

"I_1=\int u du"

"=\frac{u^2}{2}+C_1" where "C_1" = integral constant

substituting "u = tanx" back

"I_1=\frac{tan^2(x)}{2}+C"

Also "I_2=\int tan(x)dx=-ln|cos(x)|+C_2" where "C_2" = integral constant

Substituting value of "I_1" and "I_2" in equation 1

"I=\frac{tan^2(x)}{2}+ln|cos(x)|+C" where "C=C_1+C_2," another integration constant.



Another method: Using direct reduction formula (if you know or remember, otherwise please use above method).

"\int \:tan^n\left(x\right)dx=\frac{tan^{n-1}\left(x\right)}{n-1}-\int \:tan^{n-2}\left(x\right)dx+C" where "n\ne1"

Putting n =3

I "=\frac{tan^2\left(x\right)}{2}-\int \:tan\left(x\right)dx"

I "=\frac{tan^2\left(x\right)}{2}-(-ln|cos(x)|)+C"

I "=\frac{tan^2(x)}{2}+ln|cos(x)|+C"